OpenStudy (anonymous):

The ratio of two numbers is 1 to 3, and the sum of the numbers is 28. What is the smaller number?

OpenStudy (anonymous):

If their ratio is 1:3 then their sum is four times the smaller. But that's also 28, so the smaller is 7.

OpenStudy (campbell_st):

why not try x + 3x = 28 and solve for x

OpenStudy (anonymous):

(1/3) = (x/y), where x and y represent the two numbers. You can cross multiply these fractions to find that 3x = y. Dat is good info-mation! Now, we know that the sum of the two numbers is 28. So x+y=28. If 3x=y, then you can substitute in the value for y. Here: x+3x=28 Then, 4x=28, and x = 7! Y is clearly 3 times bigger than x, so the little number is 7. ^_^ Hope I helped!